Required information

E11-4 (Algo) Reporting Stockholders' Equity LO11-1, 11-2, 11-3

[The following information applies to the questions displayed below.]

The financial statements for Highland Corporation included the following selected information:

Common stock $ 1,030,000
Retained earnings $ 760,000
Net income $ 1,050,000
Shares issued 103,000
Shares outstanding 70,000
Dividends declared and paid $ 650,000

The common stock was sold at a price of $27 per share.

1. What is the amount of additional paid-in capital?

2. What was the amount of retained earnings at the beginning of the year?

3. How many shares are in treasury stock?

4. Compute earnings per share. (round to 2 decimal places)
Answer #1

Solution 1:

Amount of additional paid-in capital = (Shares issued *issue price) - Common stock

= (103000*$27) - $1030000 = $1,751,000

Solution 2:

Retained earnings at the beginning of the year = Retained earnings at the End of the year - Net income + Dividend

= $760000 - $1050000 + $650000 = $360,000

Solution 3:

Shares in treasury stock = Shares issued - Shares outstanding = 103000 - 70000 = 33,000 shares

Solution 4:

Earnings per share = Net income / shares outstanding = $1050000/ 70000 = $15 per share